OpenStudy (anonymous):

Mark deposited $9,000 into two savings accounts bearing simple interest. One of the accounts has an interest rate of 8% while the other rate is 7%. If the total interest earned after one year is $700, find the amount deposited into each of the accounts. Amount of 8% ___ Amount of 7% ___

OpenStudy (anonymous):

Can anyone HELP ME PLEASE??!!!!!!

OpenStudy (debbieg):

Suppose you let the amount in one of the accounts - let's say the 8% account, by x. Then what is an expression involving x, for the amount deposited in the 7% account?

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

Not "ok", lol... I asked a question: what is an expression involving x, for the amount in the other account?

OpenStudy (debbieg):

x goes into a 8% account How much goes into the 7 % account?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

y = 7%

OpenStudy (debbieg):

Well, we are trying to make it an expression involving x though

OpenStudy (debbieg):

Nooo, no no. 7% is the interest rate. Not an amount.

OpenStudy (debbieg):

If I deposit $9000 total, and x of that goes into 1 account, then how much goes into the other account?

OpenStudy (anonymous):

we dont know... so Y??

OpenStudy (debbieg):

You are right that we don't know. But we don't need a 2nd variable. We want an equation in terms of ONE variable, because then we can set up ONE equation.

OpenStudy (anonymous):

I really have no idea

OpenStudy (debbieg):

If I said to you: "I deposited $9000 in two accounts. I put $3000 in one." Then how much would you tell me I put in the other?

OpenStudy (anonymous):

.07(y+700)

OpenStudy (anonymous):

(9000-y)=700

OpenStudy (anonymous):

(9000-Y)

OpenStudy (anonymous):

(9000-X) (9000-Y)

OpenStudy (anonymous):

[{(9000-X)*8}/100 +{X*7}/100 ]=700 solve for X

OpenStudy (anonymous):

one moment

OpenStudy (debbieg):

@Algebra2013 do you see that, if the total amount is $9000, and you put X in one account then the AMOUNT that you put in the OTHER account has to be: $9000 - X It's important to be able to decipher the words, to understand the relationships between the quantities in the problem and set up the equation.

OpenStudy (anonymous):

Does this look right? \[\frac{ 72000-8x }{ 100 }+\frac{ 7x }{ 100 }\]

OpenStudy (debbieg):

I would not use fractions here. IMO when working with interest rates, decimals are easier and more intuitive.

OpenStudy (debbieg):

Look, he put x into the 7% account... so the interest earned from THAT ACCOUNT is: 0.07x

OpenStudy (debbieg):

the interest rate * the amount deposited. Do you understand that?

OpenStudy (anonymous):

so it would be 72000-.08X+.07X=700 ??

OpenStudy (debbieg):

No. I have no idea where you are getting the 72000 from. Please - take one step at at time. Do you understand my post above?

OpenStudy (anonymous):

.08X + .07Y

OpenStudy (anonymous):

u r getting it right solve for X and the hence the answer

OpenStudy (anonymous):

you will get X=2000

OpenStudy (anonymous):

in one account =7000 and the other X=2000

OpenStudy (anonymous):

which is 8%

OpenStudy (debbieg):

The equations are NOT right. He does not understand how to DO this problem. I'm trying to teach him something, which is the idea of this site, not to just hand out answers.

OpenStudy (anonymous):

which equation? How would you do it?

OpenStudy (anonymous):

which account has $7000 and which has $2000? 8%___ 7%___

OpenStudy (anonymous):

(7000*8)/100 +(2000*7)/100 =?

OpenStudy (anonymous):

got it 560 + 140 = 700 THank YOU

OpenStudy (anonymous):

you are welcome :)